Q1: A,B and C can do a piece of work in 36,24 and 18 days respectively.if A is assisted by B on one day and by C on next day alternatvely, how soon can work be done?
Q2: what is the compound interest in 2 and half years at 8% on $ 2500?
Q3:on a certain sum of money the compound interest for 3 years is 480 and simple interest is 460.find the rate of interest and sum of money?

Answer by stanbon(75887)

Q1: A,B and C can do a piece of work in 36,24 and 18 days respectively.if A is assisted by B on one day and by C on next day alternatvely, how soon can work be done?
A's rate = 1/36 job/day
B's rate = 1/24 job/day
C's rate = 1/18 job/day
--------------------------------
EQUATION:
Work completed on 1st day: (1/36) + (1/24) = (2+3)/72 = 5/72
Work completed on 2nd day: (1/36) + (1/18) = (3/36) = 6/72
-------------
two day total = 11/72 of the job
1st 6 days : 6(11/72) = 66/72 of the job is completed
Day 7: A and B complete 5/72 of the job
Day 8: A and C complete the job.
==================================
Q2: what is the compound interest in 2 and half years at 8% on $ 2500?
A(2.5) = 2500(1+0.08)^2.5
A(2.5) = $3030.40
--------------------------------
Q3:on a certain sum of money the compound interest for 3 years is 480 and simple interest is 460.find the rate of interest and sum of money?
Two equations:
Simple interest: Pr3 = 460
Compound interest: P[(1+r)^3 - 1) = 480
----
Solve for r and for P
============================
Cheers,
Stan H.